21. A lot consists of 144 ball pens of which 20 are defective and the others are good. Nuri will buy a pen if it is good, but will not buy if it is defective. The shopkeeper draws one pen at random and gives it to her. What is the probability that

(i) She will buy it?

(ii) She will not buy it?

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
The total numbers of outcomes i.e. pens = 144 Given, numbers of defective pens = 20 ∴ The numbers of non defective pens = 144-20 = 124 P(E) = (Number of favourable outcomes/ Total number of outcomes) (i) Total numbers events in which she will buy them = 124 So, P (buying) = 124/144 = 31/36 = 0.86 (ii) Total numbers events in which she will not buy them = 20 So, P (not buying) = 20/144 = 5/36 = 0.138